OpenStudy (anonymous):

log(x+2)-logx=1 find x

terenzreignz (terenzreignz):

Oh logs... what base are we using?

OpenStudy (tanya123):

\(log(x)+1 = log(x+2)\)

OpenStudy (dan815):

log rules

OpenStudy (dan815):

log(x+2)-logx = log((x+2)/x)=1

OpenStudy (raden):

1 is log 10 :v

OpenStudy (dan815):